# Scalable Bayesian Inference for Bradley–Terry Models with Ties: An Application to Honour Based Abuse

### v1

@@ 3.2 Markov Chain Monte Carlo
In order to sample from the posterior distribution, we develop a bespoke MCMC algorithm, which is shown in Algorithm 1. Model (1) is invariant to translations and hence an identifiability constraint is necessary. To address this, at each iteration of the MCMC algorithm, we translate the ward qualities. This does not change the value of the likelihood function as it is invariant to translations. We adapt the translation proposed in Caron and Doucet, (2012) and define
@@
4 Simulation Study
To show the effectiveness and scalability of our method, we ran a simulation study. We generated 25 synthetic data sets, including ties, on FGM abuse in South Yorkshire and sought to infer the abuse parameters.
@@ 4 Simulation Study
For each data set we draw a set of abuse parameters $\boldsymbol{\lambda}$ from the prior distribution, simulate a set of comparisons and fit the model to comparisons infer the abuse parameters. To construct the prior distribution, we incorporate spatial information, as proposed in Seymour et al., 2022b , by assuming that the levels of abuse in nearby wards had higher correlation than wards than are far away fro each other. We construct a network representation of South Yorkshire, treating each ward as a node and placing edges between adjacent wards. To describe the correlation in terms of the network, we use the matrix exponential of the network adjacency matrix, $A$. This assigns high prior correlation to pairs of wards that are highly connected and low correlation to pairs of wards that are not. We construct the covariance matrix, by computing
$\Sigma=\alpha^{2}D^{-1/2}e^{A}D^{-1/2},$
@@
where $\alpha^{2}$ is a variance parameter and $D$ is a diagonal matrix containing the elements on the main diagonal of $e^{A}$. Multiplying the matrix exponential on both sides by $D^{-1/2}$ has the effect of normalising the matrix, so we can describe the prior variance through $\alpha^{2}$. Thus the prior distribution is $\boldsymbol{\lambda}\sim N(\boldsymbol{0},\Sigma)$.
For each synthetic data set, we simulated 800 comparisons according to the model in Equations 3 and 4. This is inline with the number of comparisons we collected in South Yorkshire. We set the tie parameter $\theta\sim U[0,1]$. We fitted the model using the Pólya-Gamma latent variable representation, the Metropolis-Hasting Random Walk algorithm proposed in Seymour et al., 2022b . For the Pólya-Gamma latent variable representation, we run the MCMC algorithm for 5,000 iterations removing the fist 100 iterations as a burn-in period. For the Metropolis-Hastings Random Walk algorithm, we ran the MCMC for 500,000 iterations removing the first 100,000. iterations as a burn-in period.
For each of the simulated data sets, we recorded the effective sample size per second for the tie parameter $\delta$ and the average effective sample size per second across all risks parameters $\boldsymbol{\lambda}$. Efficient samplers will have a higher effective sample size per second. Histograms of these values are shown in Figure 2. For the Metropolis-Hastings Random Walk representation, the average effective sample size per second for the risk parameters $\boldsymbol{\lambda}$ across all simulations was 0.035 (0.070 in the most efficient simulation and 0.006 in the least). The Polya-Gamma latent variable representation offers a marked improvement, with an average effective sample size per second for the risk parameters of 108 (121 in the most effective simulation, and 91.1 in the least). As both representations use a Metropolis-Hastings Random Walk algorithm to sample values for the tie parameter $\delta$, we saw a smaller but still considerable difference. The average effective sample size per second across all simulations for $\delta$ was 13.6 (34.4 in the most effective simulation, and 2.60 in the least) using the Metropolis-Hastings Random Walk representation and 19.1 (29.2 in the most effective simulation, and 5.77 in the least).
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 2: Histograms showing the effective sample size per second for the simulation study. The top row shows the values for the Polya-Gamma latent variable representation and the bottom row the for Metropolis-Hasting Random Walk representation.
@@
Overall, the simulation study shows the main advantage of the Pólya-Gamma latent variable representation, which is a large improvement in fitting the Bradley–Terry model with ties to data, while allowing for correlated prior distributions on the parameters $\boldsymbol{\lambda}$. All simulations were carried out on a 2021 iMac with an M1 chip.
@@ 5 Results
5.1 FGM in South Yorkshire
We collected 877 comparisons from 18 participants over a three week period in November and December 2022. There were 122 tied comparisons, that is the participants said the risk of FGM in these pairs of subwards was equal. We fitted the model to the data and ran the the MCMC algorithm for 10,000 iterations treating the first 100 iterations as a burn-in period. The posterior median estimates for risk of FGM in each ward and the values of the posterior variance are shown in Figure 3.
Refer to captionRefer to caption
@@ 5.1 FGM in South Yorkshire
Our estimates show that the city of Sheffield, in the South West, has the highest risk of FGM in South Yorkshire. The highest risk wards are clustered around the centre of the city, with the four highest risk wards in the county all being in the city. In the borough of Doncaster (the eastern most borough), the risk is concentrated in Town ward, with the rest of the wards in the borough having medium to low risk of FGM. In Barnsley (the north-western borough), the risk of FGM is predominantly to the east of the borough, with the northern wards having the lowest risk in the whole of South Yorkshire. The uncertainty in our estimates is highest in Barnsley, as this was the ward where we recruited the fewest participants.
The posterior median value for the tie parameter $\delta$ is 0.468 (95% CI (0.390, 0.552)). Using the posterior median value for $\delta$ means that if a judge is shown a pair of wards with identical risks of FGM, the probability of a the judge declaring a tie is 0.139. This suggests the effect of ties were important in this study, not only reducing the risk of survey fatigue, but also providing us with data about risk of FGM in the city.
Refer to captionRefer to caption
@@ 5.2 Honour Based Abuse in Oxfordshire
We collected 766 comparisons from 12 participants in Autumn 2023. 199 of the comparisons were tied, representing 26.0% of the data set. We fitted the model to the data and ran the the MCMC algorithm for 10,000 iterations treating the first 100 iterations as a burn-in period. The posterior median estimates for risk of FGM in each ward and the values of the posterior variance are shown in Figure 5. Oxford shows a clear spatial trend, with wards in the south east of the city having the highest risk of HBA. In Banbury, the vast majority of wards are estimated to have parameter values near 0, which is the average across all wards in Oxford and Banbury. There are 9 higher risk wards on the eastern and western outskirts of the town. One ward in Banbury, East Grimsbury, is estimated to have the highest risk of HBA out of all the wards in Oxford and Banbury. The parameter value for East Grmisbury is particularly extreme. This is because the ward only featured in three comparisons, and each time it was chosen as the higher risk ward. This means the model assigned high uncertainty to this parameter. Had we used independent prior distributions, the estimated would have been more extreme. One advantage of Overall, we received more comparisons about wards in Oxford than Banbury, leading to lower uncertainty in our estimates for Oxford.
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 5: Top: Posterior median estimates and variance for the relative risk of HBA in Oxford. Bottom: Posterior median estimates and variance for the relative risk of HBA in Banbury.
We can also investigate the behaviours associated with tied comparisons. The posterior median estimate for $\delta$ is 0.964. Figure 6 shows the probability of a tie against the difference in values HBA. For pairs of ward with identical levels of HBA, we estimate that a judge would say that the wards are tied with probability 0.193.
Refer to captionRefer to caption
@@ 6 Conclusion
We used our method on two new data sets about Honour Based Abuse in UK counties, providing community level information this crime for the first time in the counties. Allowing for tied comparisons, allowed the judges to provide more accurate information about the risk of this crime. We designed a bespoke interface to collect this data and allow for tied judgement to be made. We collected over 750 comparisons in each county. Our results being used to inform community based interventions in the counties. For each study, we produced a two-page non-technical briefing describing the results (see Supplementary Material), and shared these with our project partners, local councillors and Members of the UK Parliament. Two questions were tabled in the House of Commons about Honour Based Abuse based as a result of this work – one to the Secretary of State for Health and Social Care HC Deb, 2023b and one to the Home Secretary HC Deb, 2023a .
@@ Acknowledgements
We thank South Yorkshire Police and Oxford Against Cutting for their support with data collection. Digital Research Support was provided by the Birmingham Environment for Academic Research.
@@
The data we collected and routines to implement the models and methods are available in the speedyBBT package at https://github.com/rowlandseymour/speedyBBT.

### v2

@@ 3.1 Markov chain Monte Carlo
In order to sample from the posterior distribution, we develop a bespoke MCMC algorithm, which is shown in Algorithm 1. Model (1) is invariant to translations and hence an identifiability constraint is necessary. To address this, at each iteration of the MCMC algorithm, we translate the ward qualities. This does not change the value of the likelihood function as it is invariant to translations. We adapt the translation proposed in [3] and define
@@
We have developed an R package that implements our MCMC algorithm. The speedyBBT R package is available at https://github.com/rowlandseymour/speedyBBT.
4 Simulation studies
To show the effectiveness and scalability of our model and latent variable MCMC algorithm, we ran several simulation study and sensitivity analyses. All simulations were carried out on a 2021 iMac with an M1 chip.
4.1 Testing efficiency
@@
To compare our latent variable MCMC algorithm’s efficiency to efficiently fit the BT model with ties to data against the Metropolis-Hastings random walk algorithm, we carried out a simulation study. We generated 25 synthetic data sets, including ties, on FGM abuse in South Yorkshire and sought to infer the abuse parameters using both MCMC algorithms. For each data set we draw a set of abuse parameters $\boldsymbol{\lambda}$ from the prior distribution, simulate a set of comparisons and fit the model to comparisons infer the ward quality parameters. To construct the prior distribution, we incorporate spatial information, as proposed in [16], by assuming that the levels of abuse in nearby wards had higher correlation than wards that are far away from each other. We construct a network representation of South Yorkshire, treating each ward as a node and placing edges between adjacent wards. To describe the correlation in terms of the network, we use the matrix exponential of the network adjacency matrix, $A$. This assigns high prior correlation to pairs of wards that are highly connected and low correlation to pairs of wards that are weakly connected. We construct the covariance matrix, by computing
$\Sigma=\alpha^{2}D^{-1/2}e^{A}D^{-1/2},$ (12)
@@
where $\alpha^{2}$ is a signal variance parameter describing the overall strength of correlation in the prior covariance matrix, and $D$ is a diagonal matrix containing the elements on the main diagonal of $e^{A}$. Multiplying the matrix exponential on both sides by $D^{-1/2}$ has the effect of normalising the matrix, so we can describe the prior variance through $\alpha^{2}$. Small values of $\alpha^{2}$ correspond to applications with very strong spatial correlation and as $\alpha^{2}$ increase the prior correlation decreases in strength. The prior distribution is $\boldsymbol{\lambda}\sim N(\boldsymbol{0},\Sigma)$.
For each synthetic data set, we simulated 800 comparisons according to the model in Equations 4 and 6. This is inline with the number of comparisons we collected in South Yorkshire. We draw a value of the tie parameter $\delta\sim U[0,1]$. We set the prior variance parameter $\alpha^{2}=1$.
@@
For each of the simulated data sets, we recorded the effective sample size per second for the tie parameter $\delta$ and the average effective sample size per second across all risks parameters $\boldsymbol{\lambda}$. Efficient samplers will have a higher effective sample size per second. Histograms of these values are shown in Figure 2. Overall, the simulation study shows the main advantage of the Pólya-Gamma latent variable representation, which is a large improvement in fitting the Bradley–Terry model with ties to data, while allowing for correlated prior distributions on the parameters $\boldsymbol{\lambda}$. Further details can be found in the Supplementary Material
Figure 2: Histograms showing the effective sample size per second for the simulation study. The top row shows the values for the Pólya-Gamma latent variable representation and the bottom row the for Metropolis-Hastings random walk representation.
4.2 Testing scalability
To test the scalability of our MCMC algorithm, we ran a simulation study comparing it to a Metropolis-Hastings random walk algorithm. We generated synthetic data sets with 16, 32, 64, 128, 256, 512, and 1024 wards. For each data set we simulated a prior covariance matrix from a Wishart distribution with the identity matrix as the scale matrix and the number of wards as the degrees of freedom. We then normalised that matrix as shown in Equation (12). We drew a set of ward quality parameters from a multivariate normal distribution with zero mean and covariance matrix given by the normalised synthetic prior covariance matrix. Finally, we generated a synthetic set of comparisons using the simulated ward quality parameters. The number of comparisons we generated was ten times the number of wards, as this is the recommended amount of data to collected in a comparative judgement study [11] and close to the amount we collected in South Yorkshire. We also fixed $\delta=0.5$ when simulating the comparisons as this is similar to the value we inferred in the South Yorkshire data set.
For each value of the number of wards, we generated ten synthetic data sets and fitted the model to each one using the latent variable algorithm and the Metropolis-Hastings algorithm described in [16]. We ran our latent variable algorithm for 5,000 iterations, treating the first 100 iterations as a burn-in period. The Metropolis-Hastings random walk algorithm took longer to converge, and we ran it for 100,000 iterations removing the first 1,000 as a burn-in period. We did not fit the Metropolis-Hastings random walk to the data sets with 1,024 wards as we expected each of the ten runs to take around 6 days. Figure 3 shows the mean time taken to run the MCMC algorithms across each of the 10 runs. Our latent variable representation provides significant time advantages compared to the Metropolis-Hastings random walk algorithm and scales well as the number of wards is increased. With 1,024, our latent variable representation took on average 20 minutes to run. For data sets of the size we collected with our partners (64 - 128 wards), the model took around 20 to 40 seconds to run. This compares with 8 to 77 minutes using the Metropolis-Hastings random walk algorithm.
Figure 3: The mean time take to fit the the models to data sets of different sizes. The clear points correspond to the latent variable algorithm and the solid points to the Metropolis-Hastings random walk algorithm. Both axes are shown on a log scale.
4.3 Sensitivity to the tie parameter and prior variance hyperparameter
To determine if our model was sensitive to the value of the prior variance parameter $\alpha^{2}$ in Equation 12 or the value of the tie parameter $\delta$, we carried out a simulation study for each parameter. To determine the sensitivity to the prior variance parameter, we fitted the model to the South Yorkshire data fixing the value of $\alpha^{2}$ to different values. These values represent, very strong, strong, and weak levels of prior spatial correlation, and the values were chosen based on the end points of the 95% credible interval. We found that there was little difference in the estimates for the the ward quality parameters when using strong or weak prior spatial correlation to when inferring the parameter. Only when assuming very strong prior correlation (fixing $\alpha^{2}$ to a small value outside of the 95% credible interval when it is learned) did we see a difference. In this simulation, there was substantial spatial smoothing and shrinkage of the ward quality parameters. Figures showing the effect of this smoothing are shown the Supplementary Material.
To assess the sensitivity of our model to the tie parameter $\delta$, we simulated different data sets of FGM in South Yorkshire using different values of $\delta$. The values were chosen such that approximately 5%, 20%, 50% and 75% of the comparisons were tied. We found we could accurately recover the value of the tie and ward quality parameters, however there was a slight increase in the error when the data set comprises of 5% and 75% of ties. In the case where 5% of the comparisons were tied, the model had less information to learn about the value of $\delta$ and made small adjustments to the values of the ward quality parameters to account for this. In contrast, when 75% of the comparisons were tied, the model lacked enough information to accurately infer the values of the ward quality parameters. Given in the data we collected, around 15% and 25% of the comparisons are tied, we believe we are able to fully recover all model parameters.
Full details of both of these analyses are presented in the Supplementary Material.
@@ 5 Results
5.1 Female genital mutilation in South Yorkshire
We collected 877 comparisons from 18 participants over a three week period in November and December 2022. There were 122 tied comparisons, that is the participants said the risk of FGM in these pairs of subwards was equal. We fitted the model to the data and ran the the MCMC algorithm for 5,000 iterations treating the first 100 iterations as a burn-in period. The prior distribution with spatial correlation for the ward quality parameters was given by network structure, as described in Section 4. For this prior distribution’s variance parameter $\alpha^{2}$, we use a conjugate inverse-gamma prior distribution with shape and scale set to 0.01. For the tie parameter $\delta$, we use an Exp(0.01) prior distribution, as this is sufficiently vague. The posterior median estimates for risk of FGM in each ward and the values of the posterior variance are shown in Figure 4. Trace plots and convergence diagnoses can be found in the Supplementary Material.
@@ 5.1 Female genital mutilation in South Yorkshire
Our estimates show that the city of Sheffield, in the South West, has the highest risk of FGM in South Yorkshire. The highest risk wards are clustered around the centre of the city, with the four highest risk wards in the county all being in the city. In the borough of Doncaster (the eastern most borough), the risk is concentrated in Town ward, with the rest of the wards in the borough having medium to low risk of FGM. In Barnsley (the north-western borough), the risk of FGM is predominantly to the east of the borough, with the northern wards having the lowest risk in the whole of South Yorkshire. The uncertainty in our estimates is highest in Barnsley, as this was the borough where we recruited the fewest participants.
The posterior median value for the tie parameter $\delta$ is 0.468 (95% CI (0.390, 0.552)). Using the posterior median value for $\delta$ means that if a judge is shown a pair of wards with identical risks of FGM, the probability of a the judge declaring a tie is 0.139. This suggests the effect of ties were important in this study, not only reducing the risk of survey fatigue, but also providing us with information about risk of FGM in the city.
@@ 5.2 Honour based abuse in Oxfordshire
We collected 766 comparisons from 12 participants in Autumn 2023. 199 of the comparisons were tied, representing 26.0% of the data set. We fitted the model to the data and ran the MCMC algorithm for 5,000 iterations treating the first 100 iterations as a burn-in period. Trace plots and convergence diagnoses can be found in the Supplementary Material.
The posterior median estimates for risk of FGM in each ward and the values of the posterior variance are shown in Figure 6. Oxford shows a clear spatial trend, with wards in the south east of the city having the highest risk of honour based abuse. In Banbury, the vast majority of wards are estimated to have parameter values near 0, which is the average across all wards in Oxford and Banbury. There are 9 higher risk wards on the eastern and western outskirts of the town. One ward in Banbury, East Grimsbury, is estimated to have the highest risk of honour based abuse out of all the wards in Oxford and Banbury. The parameter value for East Grmisbury is particularly extreme. This is because the ward only featured in three comparisons, and each time it was chosen as the higher risk ward. This means the model assigned high uncertainty to this parameter. Had we used independent prior distributions, the estimated would have been more extreme. Overall, we received more comparisons about wards in Oxford than Banbury, leading to lower uncertainty in our estimates for Oxford.
Figure 6: Top: Posterior median estimates and variance for the relative risk of honour based abuse in Oxford. Bottom: Posterior median estimates and variance for the relative risk of honour based abuse in Banbury.
We can also investigate the behaviours associated with tied comparisons. The posterior median estimate for $\delta$ is 0.964. Figure 7 shows the probability of a tie against the difference in values honour based abuse. For pairs of ward with identical levels of honour based abuse, we estimate that a judge would say that the wards are tied with probability 0.193.
@@ 6 Conclusion
We used our method on two new data sets about honour based abuse in UK counties, providing community level information this crime for the first time in the counties. Allowing for tied comparisons, allowed the judges to provide more accurate information about the risk of this crime. We designed a bespoke interface to collect this data and allow for tied judgement to be made. We collected over 750 comparisons in each county. Our results being used to inform community based interventions in the counties. For each study, we produced a two-page non-technical briefing describing the results (see Supplementary Material), and shared these with our project partners, local councillors and Members of the UK Parliament. Two questions were tabled in the House of Commons about Honour Based Abuse based as a result of this work – one to the Secretary of State for Health and Social Care [10] and one to the Home Secretary [9].
@@ Acknowledgements
We thank South Yorkshire Police and Oxford Against Cutting for their support with data collection. Digital Research Support was provided by the Birmingham Environment for Academic Research. This work was supported by a UK Research and Innovation Future Leaders Fellowship [MR/X034992/1]. We thank the two anonymous reviewers and the Associate Editor for their constructive comments.
@@
We have developed an R package to fit models to the data. The speedyBBT R package is available at CRAN [15] and also contains the data. The python interface we developed to collect the data is available at https://gitlab.bham.ac.uk/seymourg-bsbt/comparison-interface [17].